OpenStudy (anonymous):

I THINK IT'S D AM I RIGHT HELP !! Select the best possible first step to solving the system by first eliminating the x variable. 4x + 9y = 21 3x - 8y = 1 Select one: a. Multiply the first equation by 3 and multiply the second equation by 4. b. Multiply the first equation by -3 and multiply the second equation by -4. c. Multiply the first equation by 8 and multiply the second equation by 9. ***d. None of the above

OpenStudy (anonymous):

a is the answer its really easy

OpenStudy (anonymous):

BUT WHY ?

OpenStudy (jjuden):

it is A

OpenStudy (anonymous):

because if you do what a says the x variable is the same and will cancel each other out

OpenStudy (anonymous):

yea it is d cause you need them to be opposite signs to cancel them out

OpenStudy (anonymous):

one needs to be positive and one need to be negitive otherwise it can't cancel out

OpenStudy (anonymous):

no you do that when you put them into one equation you change the sign of one of the equations to negative thats the second step not first

OpenStudy (anonymous):

not how i learned it.to cancel out they do need to be the same number but they have to be opostie signs.one will be positive 12 the other is negetive 12 and when added they cancel out

OpenStudy (anonymous):

i have done algebra since I was 8 I know what I am talking about

OpenStudy (anonymous):

think man if they are both positive 12 they will ad to 24.i should know cause ive done this stuff for 4 years now

OpenStudy (anonymous):

actually 5 years cause im in 10th grade now.ive done it since 5th grade

OpenStudy (anonymous):

because to cancel out you would do this just an example 2x+4=8 2x+6=12 2x+4-(2x+4)=8-12 I showed the whole thing \

OpenStudy (anonymous):

and if the x variable was not the same you would multiply it by a number that would make them be the same

OpenStudy (anonymous):

a

OpenStudy (amistre64):

i would contend that a and b are equally best as options go

OpenStudy (anonymous):

@amistre64 true

OpenStudy (anonymous):

so it's A?

OpenStudy (anonymous):

(4x + 9y = 21)*-3=-12x-27y=-63 (3x - 8y)*4= 1*4 =12x+32y=4 12x+(-12x)=0 -27y=-63 32y=4 4y=4 y=1

OpenStudy (anonymous):

but i think they are looking for a, though

OpenStudy (anonymous):

yes, a

OpenStudy (amistre64):

my thought is that it is a poorly worded question and that they are expecting you to pick d

OpenStudy (anonymous):

that is what i said @amistre64

OpenStudy (anonymous):

what about marines work is it wrong? yeah thats what i thought and thats what marine said but everyone says it A

OpenStudy (amistre64):

none of the options actually eliminates the x parts ... even tho a and b are viable first steps to take in the process

OpenStudy (anonymous):

i would think it would be a because it is the first step

OpenStudy (anonymous):

but it is asking you to eliminate

OpenStudy (anonymous):

to eliminate you multiply them by a common variable but one ALWAYS has to be negitive to cancel them out which in this case is -3 and +4

OpenStudy (amistre64):

id just multiply the top by -3/4 :)

OpenStudy (anonymous):

When analyzing three linear functions, Michele noticed that they all intersected at the point (5, 3). What conclusion can Michele come to about the linear function with the smallest slope? Select one: a. It has the middle y-intercept. b. It has the highest y-intercept. c. It has the lowest y-intercept. d. Michele cannot conclude anything.

OpenStudy (anonymous):

what about this one its A right?

OpenStudy (anonymous):

unless it shows the functions it is hard to tell

OpenStudy (anonymous):

i think it would be d

OpenStudy (anonymous):

i have to agree with athenasannabeth on this one

OpenStudy (anonymous):

Connor went to New York for vacation. He spent 3 nights at a hotel and rented a car for 4 days. Jillian stayed at the same hotel, but spent 4 nights and rented a car for 5 days from the same company. If Connor paid $675 and Jillian paid $875, how much did one night at the hotel cost? Select one: a. $75 b. $100 c. $125 d. $150

OpenStudy (anonymous):

i said it was 100 but how can i know if im right?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

this is a tough one.im not sure

OpenStudy (anonymous):

When planning his crops, Farmer Tim knows that his 7 acres can support peas and carrots. He wants to make $7,000 from his crops. He can make $2,000 per acre of peas (the variable p) and $1,000 per acre of carrots (the variable c). Which equation below would be a constraint in his system of equations? Select one: a. 2,000p + 1,000c = 7,000 b. 2,000p - 1,000c = 7,000 c. p + c = 7 d. c - p = 7

OpenStudy (anonymous):

what about this one marines ???

OpenStudy (anonymous):

the one with the hotel is 125

OpenStudy (anonymous):

how did you get it?

OpenStudy (anonymous):

it would be a for the one about crops

OpenStudy (anonymous):

the reason is cause when you add the amount from peas and the one from carrots he will get 7,000 all together

OpenStudy (anonymous):

this one i think its B am i wrong?

OpenStudy (anonymous):

ITS D:)

OpenStudy (anonymous):

NO NO NO, ITS C...OR B...IDK

OpenStudy (anonymous):

BUT ITS DEFINETELY NOT A

OpenStudy (anonymous):

for that one it is a because the second equatioon is in slope intercept form and +1 it the y intercept

OpenStudy (anonymous):

you shouldn't do it the way I did but I tried every answer so 125 multiply by 3= 375 675-375=300 so thats how much he paid to rent a car then 300 divide by 4=75 then we figure out if it works for both of them 75 multiply by 5 = 375 875-375=500 500/4=125 but you should solve it by putting it in a equation oh and this is for the hotel one

OpenStudy (anonymous):

is this D?

OpenStudy (anonymous):

im guessing on this one.i think it is d.im not sure so i might need athena opinion

OpenStudy (anonymous):

probably

OpenStudy (anonymous):

this i think D?

OpenStudy (anonymous):

it is a cause they intersect at positive 2,3

OpenStudy (anonymous):

This is C??

OpenStudy (anonymous):

ITS C

OpenStudy (anonymous):

ITS C ITS C ITS C!!!!!:D

OpenStudy (anonymous):

yea it is c

OpenStudy (anonymous):

right?:D i have that class too!!!

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

lol graciekeyr

OpenStudy (anonymous):

i think that one is b

OpenStudy (anonymous):

it is probably b

OpenStudy (anonymous):

but is their more answers?

OpenStudy (anonymous):

is this a test or something?

OpenStudy (anonymous):

is that all you need help with?